# Towards Unpolarized GPDs from Pseudo-Distributions

### v1

@@ 8 Acknowledgments
This project was supported by the U.S. Department of Energy, Office of Science, Contract No. DE-AC05-06OR23177, under which Jefferson Science Associates, LLC operates Jefferson Lab. KO and HD were supported in part by U.S. DOE Grant #DE-FG02-04ER41302. CM is supported in part by U.S. DOE ECA #DE-SC0023047. AR acknowledges support by U.S. DOE Grant #DE-FG02-97ER41028. This work has benefited from the collaboration enabled by the Quark-Gluon Tomography (QGT) Topical Collaboration, U.S. DOE Award #DE-SC0023646. This research was funded, in part, by l’Agence Nationale de la Recherche (ANR), project ANR-23-CE31-0019. For the purpose of open access, the author has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission. Computations for this work were carried out in part on facilities of the USQCD Collaboration, which are funded by the Office of Science of the U.S. Department of Energy. This work was performed in part using computing facilities at William & Mary which were provided by contributions from the National Science Foundation (MRI grant PHY-1626177), and the Commonwealth of Virginia Equipment Trust Fund. This work used the Extreme Science and Engineering Discovery Environment (XSEDE), which is supported by National Science Foundation grant number ACI-1548562. Specifically, it used the Bridges system, which is supported by NSF award number ACI-1445606, at the Pittsburgh Supercomputing Center (PSC) Towns et al. (2014); Nystrom et al. (2015). In addition, this work used resources at NERSC, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract #DE-AC02-05CH11231, as well as resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. #DE-AC05-00OR22725. The software codes Chroma Edwards and Joo (2005), QUDA Clark et al. (2010); Babich et al. (2010), QPhiX Joó et al. (2016), and Redstar Chen et al. (2023) were used in our work. The authors acknowledge support from the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research and Office of Nuclear Physics, Scientific Discovery through Advanced Computing (SciDAC) program, and of the U.S. Department of Energy Exascale Computing Project. The authors also acknowledge the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing HPC resources, like Frontera computing system Stanzione et al. (2020) that has contributed to the research results reported within this paper. The authors acknowledge William & Mary Research Computing for providing computational resources and/or technical support that have contributed to the results reported within this paper.
@@ Appendix A Full set of Lorentz amplitudes
Figure 21: Full set of amplitudes for $\vec{p}_{f}=(2,1,0)$ and $\vec{p}_{i}=(0,1,2)$. See a discussion of the features of the plot in the text. The first row presents the real parts of $\mathcal{A}_{1}$ (left), $\mathcal{A}_{4}$ (middle), and $\mathcal{A}_{6}$ (right) and the second row their imaginary parts. The third row presents the real parts of $\mathcal{A}_{5}$ (left), $\mathcal{A}_{7}$ (middle), and $\mathcal{A}_{2,3,8}$ at once (right), and the fourth row their imaginary parts. In all cases but the grouped plot of $\mathcal{A}_{2,3,8}$, we show three extractions of the amplitudes: using a cut of $t_{s}=3$ (blue, see section 4) and an $l^{2}$ minimization (or SVD, see section 3.3), using a cut of $t_{s}=4$ and the $l^{2}$ minimization (green), or using a cut of 3 and an $l^{1}$ minimization (orange). In the last plot, we only use a cut of 3 with the $l^{2}$ minimization for clarity.
@@
  • For the five amplitudes that enter the construction of light-cone GPDs, we compare a solution for the extraction of the amplitudes from the matrix elements using either the SVD ($l^{2}$ minimization) or an $l^{1}$ minimization. We find that the $l^{1}$ minimization produces generally larger uncertainties but with central values in excellent agreement with respect to the $l^{2}$ extraction. We consider this as a validation of the robustness of the SVD method.
@@
  • $\mathcal{A}_{5}$ and $\mathcal{A}_{7}$ do not show a strong deviation from 0, a situation generally observed in the dataset. $\mathcal{A}_{5}$ contains signal of the $D$-term (absent in this non-singlet calculation), the kinematic higher-twist contamination $Y$ discussed in Radyushkin (2024), and lattice systematic errors. Therefore, it is quite satisfactory that no clear signal emerges. As for $\mathcal{A}_{7}$, it enters the construction of the light-cone GPD $E$, but is generally small in magnitude and barely different from 0.
  • The kinematic factors of the amplitudes $\mathcal{A}_{2,3,6,7,8}$ cancel in the limit $z=0$. Therefore, we cannot evaluate the amplitudes in this limit. $\mathcal{A}_{6}$ grows noticeably at small $z$. It is in fact necessary to counter-act the discontinuity of $\mathcal{A}_{4}$ between the value computed at $z=0$ (the form factor $F_{2}$), and the value at $z=1$. Adding $\nu\mathcal{A}_{6}$ to $\mathcal{A}_{4}$ to form the light-cone GPD $E$ alleviates the discontinuity [^6] The discontinuity of $\mathcal{A}_{4}$ is perceivable in the first row, second column plot of Fig. 21. Because $\nu$ is negative for $z>0$ with this choice of momenta $p_{f}$ and $p_{i}$, $\nu\mathcal{A}_{6}$ is negative at $z=1$ and alleviates the discontinuity of $\mathcal{A}_{4}$ to form the GPD $E$..

### v2

@@ 8 Acknowledgments
This project was supported by the U.S. Department of Energy, Office of Science, Contract No. DE-AC05-06OR23177, under which Jefferson Science Associates, LLC operates Jefferson Lab. KO and HD were supported in part by U.S. DOE Grant #DE-FG02-04ER41302. CM is supported in part by U.S. DOE ECA #DE-SC0023047. AR acknowledges support by U.S. DOE Grant #DE-FG02-97ER41028. The work of HD and DR was conducted in part under the Laboratory Directed Research and Development Program (LDRD 2412) at Thomas Jefferson National Accelerator Facility for the U.S. Department of Energy. This work has benefited from the collaboration enabled by the Quark-Gluon Tomography (QGT) Topical Collaboration, U.S. DOE Award #DE-SC0023646. This research was funded, in part, by l’Agence Nationale de la Recherche (ANR), project ANR-23-CE31-0019. For the purpose of open access, the authors have applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission. Computations for this work were carried out in part on facilities of the USQCD Collaboration, which are funded by the Office of Science of the U.S. Department of Energy. This work was performed in part using computing facilities at William & Mary which were provided by contributions from the National Science Foundation (MRI grant PHY-1626177), and the Commonwealth of Virginia Equipment Trust Fund. This work used the Extreme Science and Engineering Discovery Environment (XSEDE), which is supported by National Science Foundation grant number ACI-1548562. Specifically, it used the Bridges system, which is supported by NSF award number ACI-1445606, at the Pittsburgh Supercomputing Center (PSC) Towns et al. (2014); Nystrom et al. (2015). In addition, this work used resources at NERSC, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract #DE-AC02-05CH11231, as well as resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. #DE-AC05-00OR22725. The software codes Chroma Edwards and Joo (2005), QUDA Clark et al. (2010); Babich et al. (2010), QPhiX Joó et al. (2016), and Redstar Chen et al. (2023) were used in our work. The authors acknowledge support from the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research and Office of Nuclear Physics, Scientific Discovery through Advanced Computing (SciDAC) program, and of the U.S. Department of Energy Exascale Computing Project (ECP). The authors also acknowledge the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing HPC resources, like Frontera computing system Stanzione et al. (2020) that has contributed to the research results reported within this paper. The authors acknowledge William & Mary Research Computing for providing computational resources and/or technical support that have contributed to the results reported within this paper.
@@ Appendix A Full set of Lorentz amplitudes
Figure 25: Full set of amplitudes for $\vec{p}_{f}=(2,1,0)$ and $\vec{p}_{i}=(0,1,2)$. See a discussion of the features of the plot in the text. The first row presents the real parts of $\mathcal{A}_{1}$ (left), $\mathcal{A}_{4}$ (middle), and $\mathcal{A}_{6}$ (right), and the second row their imaginary parts. The third row presents the real parts of $\mathcal{A}_{5}$ (left), $\mathcal{A}_{7}$ (middle), and $\mathcal{A}_{2,3,8}$ at once (right), and the fourth row their imaginary parts. In all cases but the grouped plot of $\mathcal{A}_{2,3,8}$, we show three extractions of the amplitudes: using a cut of $t_{s}=3$ (blue, see section 4) and an $l^{2}$ minimization (or SVD, see section 3.3), using a cut of $t_{s}=4$ and the $l^{2}$ minimization (green), or using a cut of 3 and an $l^{1}$ minimization (orange). In the last plot, we only use a cut of 3 with the $l^{2}$ minimization for clarity.
@@
  • For the five amplitudes that enter the construction of light-cone GPDs, we compare a solution for the extraction of the amplitudes from the matrix elements using either the SVD ($l^{2}$ minimization) or an $l^{1}$ minimization. We find that the $l^{1}$ minimization produces generally larger uncertainties but with central values in excellent agreement with the $l^{2}$ extraction. We consider this as a validation of the robustness of the SVD method.
@@
  • $\mathcal{A}_{5}$ and $\mathcal{A}_{7}$ do not show a strong deviation from 0, a situation generally observed in the dataset. $\mathcal{A}_{5}$ contains signal of the $D$-term, the kinematic higher-twist contamination $Y$ discussed in Radyushkin (2024), and lattice systematic errors. As for $\mathcal{A}_{7}$, it enters the construction of the light-cone GPD $E$, but is generally small in magnitude and barely different from 0.
  • The kinematic factors of the amplitudes $\mathcal{A}_{2,3,6,7,8}$ cancel in the limit $z=0$. Therefore, we cannot evaluate the amplitudes in this limit. $\mathcal{A}_{6}$ grows noticeably at small $z$. It is in fact necessary to counter-act the discontinuity of $\mathcal{A}_{4}$ between the value computed at $z=0$ (the form factor $F_{2}$), and the value at $z=1$. Adding $\nu\mathcal{A}_{6}$ to $\mathcal{A}_{4}$ to form the light-cone GPD $E$ alleviates the discontinuity, as we demonstrate in Figure 26.
Figure 26: Effect of adding $\nu\mathcal{A}_{6}$ to $\mathcal{A}_{4}$ for $\vec{p}_{f}=(2,1,0)$ and $\vec{p}_{i}=(0,1,2)$.
@@ Appendix B To $\gamma_{3}$ or not to $\gamma_{3}$
Appendix C Effect of the range of separations $z$
We compare in Fig. 30 the extraction of the moment $A_{3,0}$ in one bin in $t$ depending on the range in $z$ one chooses to consider. The standard analysis in the main body of the article is performed with $z\in[a,6a]$. We observe that varying this range gives statistically similar results, not least because of the very high correlation between the values in $z$. The very inaccurate result at small $z$ stems from the fact that $A_{3,0}$ is a subdominant moment in the real part of the Ioffe-time distribution, which can only be captured accurately when the Ioffe time is large enough.
Figure 30: The red points are the matched $A_{3,0}(t=-0.33\textrm{ GeV}^{2},\mu=2$ GeV) depending on the non-local separation $z$ for a cut $t_{s}=4$. The bands represent the correlated fit by a constant of a subset of the red points. It appears that the correlated fit result is very stable whereas an uncorrelated fit would show a significant variation. We observed a similar robustness of the correlated constant fits for the non-local extraction of the EFF.
Figure 31: $A_{2,0}(t,\mu=2$ GeV) (left) and $B_{2,0}(t,\mu=2$ GeV) (right) extracted from the lattice cut of 3 by fitting the matched values for $z\in[a,6a]$ (red) or $z\in[a,4a]$.
Figure 32: $A_{3,0}(t,\mu=2$ GeV) (left) and $B_{3,0}(t,\mu=2$ GeV) (right) extracted from the lattice cut of 3 by fitting the matched values for $z\in[a,6a]$ (red) or $z\in[a,4a]$.
In Figs. 31 to 36, we show systematically the extraction of the moments using $z\in[a,6a]$ and $z\in[a,4a]$. It appears that there is no significant statistical difference between using the two methods, and no clear sign of higher twist contributions within the precision of this study.
Figure 33: $A_{4,0}(t,\mu=2$ GeV) (left) and $B_{4,0}(t,\mu=2$ GeV) (right) extracted from the lattice cut of 3 by fitting the matched values for $z\in[a,6a]$ (red) or $z\in[a,4a]$.
Figure 34: $A_{3,2}(t,\mu=2$ GeV) (left) and $B_{3,2}(t,\mu=2$ GeV) (right) extracted from the lattice cut of 3 by fitting the matched values for $z\in[a,6a]$ (red) or $z\in[a,4a]$.
Figure 35: $A_{4,2}(t,\mu=2$ GeV) (left) and $B_{4,2}(t,\mu=2$ GeV) (right) extracted from the lattice cut of 3 by fitting the matched values for $z\in[a,6a]$ (red) or $z\in[a,4a]$.
Figure 36: $C_{2}(t,\mu=2$ GeV) extracted from the lattice cut of 3 by fitting the matched values for $z\in[a,6a]$ (red) or $z\in[a,4a]$.